# TransPixar: Advancing Text-to-Video Generation with Transparency

### v1

@@ Abstract
Text-to-video generative models have made significant strides, enabling diverse applications in entertainment, advertising, and education. However, generating RGBA video, which includes alpha channels for transparency, remains a challenge due to limited datasets and the difficulty of adapting existing models. Alpha channels are crucial for visual effects (VFX), allowing transparent elements like smoke and reflections to blend seamlessly into scenes. We introduce TransPixar, a method to extend pretrained video models for RGBA generation while retaining the original RGB capabilities. TransPixar leverages a diffusion transformer (DiT) architecture, incorporating alpha-specific tokens and using LoRA-based fine-tuning to jointly generate RGB and alpha channels with high consistency. By optimizing attention mechanisms, TransPixar preserves the strengths of the original RGB model and achieves strong alignment between RGB and alpha channels despite limited training data. Our approach effectively generates diverse and consistent RGBA videos, advancing the possibilities for VFX and interactive content creation. The code is available at https://wileewang.github.io/TransPixar/.
@@
Figure 1: RGBA Video Generation with TransPixar. By introducing LoRA layers into DiT-based text-to-video model with a novel alpha channel adaptive attention mechanism, our method enables RGBA video generation from text while preserving Text-to-Video quality.
@@ 1 Introduction
In this work, we propose TransPixar, which effectively adapts the pretrained RGB video models to generate RGB channels and the alpha channel simultaneously. We leverage state-of-the-art DiT-like video generation models [57, 26] , and additionally introduce new tokens appended after text and RGB tokens for generating the alpha channels. To facilitate convergence, we reinitialize the positional embeddings for the alpha tokens and introduce a zero-initialized, learnable domain embedding to distinguish alpha tokens from RGB tokens. Furthermore, we employ a LoRA-based fine-tuning scheme [23], applied exclusively to project alpha tokens into the qkv space, thereby maintaining RGB generation quality. With the proposed approach, we extend the modality while preserving the original input-output structure and relying on the existing attention mechanism through LoRA adaptation.
@@ 2 Related Work
Video Matting. A straightforward approach for RGBA video generation is to extract the alpha channel from generated RGB content, as done with traditional green screen keying or learning-based video matting expert models [28, 29, 30]. OmnimatteRF [28] introduces a video matting method that combines dynamic 2D foreground layers with a 3D background model, enabling more realistic scene reconstruction for real-world videos. Robust Video Matting (RVM) [30] proposes a real-time, high-quality human video matting method with a recurrent architecture for improved temporal coherence, achieving state-of-the-art results without auxiliary inputs. Another work presents a high-speed, high-resolution background replacement technique with precise alpha matte extraction, supported by the VideoMatte240K and PhotoMatte13K/85 datasets [29]. Additionally, many image matting methods [5, 27, 58, 51] can be applied for frame-by-frame matting.
@@ 3 Method
Figure 3: Pipeline of TransPixar. Our method is organized as follows: (1) Left: we extend the input of DiT to include new alpha tokens; (2) Top Center: we initialize alpha tokens with our positional encoding; (3) Bottom Center: we insert a partial LoRA and adjust attention computation during training and inference.

### v2

# TransPixeler: Advancing Text-to-Video Generation with Transparency

@@ Abstract
Text-to-video generative models have made significant strides, enabling diverse applications in entertainment, advertising, and education. However, generating RGBA video, which includes alpha channels for transparency, remains a challenge due to limited datasets and the difficulty of adapting existing models. Alpha channels are crucial for visual effects (VFX), allowing transparent elements like smoke and reflections to blend seamlessly into scenes. We introduce TransPixeler, a method to extend pretrained video models for RGBA generation while retaining the original RGB capabilities. TransPixeler leverages a diffusion transformer (DiT) architecture, incorporating alpha-specific tokens and using LoRA-based fine-tuning to jointly generate RGB and alpha channels with high consistency. By optimizing attention mechanisms, TransPixeler preserves the strengths of the original RGB model and achieves strong alignment between RGB and alpha channels despite limited training data. Our approach effectively generates diverse and consistent RGBA videos, advancing the possibilities for VFX and interactive content creation. The code is available at https://wileewang.github.io/TransPixeler/.
@@
Figure 1: RGBA Video Generation with TransPixeler. By introducing LoRA layers into DiT-based text-to-video model with a novel alpha channel adaptive attention mechanism, our method enables RGBA video generation from text while preserving Text-to-Video quality.
@@ 1 Introduction
In this work, we propose TransPixeler, which effectively adapts the pretrained RGB video models to generate RGB channels and the alpha channel simultaneously. We leverage state-of-the-art DiT-like video generation models [58, 27] , and additionally introduce new tokens appended after text and RGB tokens for generating the alpha channels. To facilitate convergence, we reinitialize the positional embeddings for the alpha tokens and introduce a zero-initialized, learnable domain embedding to distinguish alpha tokens from RGB tokens. Furthermore, we employ a LoRA-based fine-tuning scheme [24], applied exclusively to project alpha tokens into the qkv space, thereby maintaining RGB generation quality. With the proposed approach, we extend the modality while preserving the original input-output structure and relying on the existing attention mechanism through LoRA adaptation.
@@ 2 Related Work
Video Matting. A straightforward approach for RGBA video generation is to extract the alpha channel from generated RGB content, as done with traditional green screen keying or learning-based video matting expert models [29, 30, 31]. OmnimatteRF [29] introduces a video matting method that combines dynamic 2D foreground layers with a 3D background model, enabling more realistic scene reconstruction for real-world videos. Robust Video Matting (RVM) [31] proposes a real-time, high-quality human video matting method with a recurrent architecture for improved temporal coherence, achieving state-of-the-art results without auxiliary inputs. Another work presents a high-speed, high-resolution background replacement technique with precise alpha matte extraction, supported by the VideoMatte240K and PhotoMatte13K/85 datasets [30]. Additionally, many image matting methods [6, 28, 59, 52, 4] can be applied for frame-by-frame matting.
@@ 3 Method
Figure 3: Pipeline of TransPixeler. Our method is organized as follows: (1) Left: we extend the input of DiT to include new alpha tokens; (2) Top Center: we initialize alpha tokens with our positional encoding; (3) Bottom Center: we insert a partial LoRA and adjust attention computation during training and inference.